# Quantum integration of decay rates at second order in perturbation theory

### v1

@@ V Conclusions
We have presented the first quantum computation of a total decay rate at second order in perturbative quantum field theory. Leveraging the loop-tree duality (LTD) framework, we have successfully combined loop and tree-level Feynman diagrams with a quantum algorithm on a quantum computer. This methodological advancement is significant from the high-energy physics perspective, as it allows us to integrate a real process with potential for quantum speedup. While we do not claim to have achieved quantum speedup in this work, our results lay the groundwork for future explorations in this direction. From the perspective of quantum computing, our study marks a noteworthy achievement. By solving a relatively complicated regression problem using a Quantum Neural Network (QNN) on a realistic dataset, we found a good compromise between trainability and expressibility, a common challenge in quantum neural networks. Most of the results presented are in agreement with the expected values within the uncertainty bands. This demonstrates the potential of quantum computing to address complex, real-world problems and highlights the importance of continuing to push the boundaries of what quantum technology can achieve.
@@ Acknowledgements.
This work is supported by the Spanish Government - Agencia Estatal de Investigación (MCIN/AEI/10.13039/501100011033) Grant No. PID2020-114473GB-I00, and Generalitat Valenciana Grants No. PROMETEO/2021/071 and ASFAE/2022/009 (Planes Complementarios de I+D+i, NextGenerationEU). This work is also supported by the Ministry of Economic Affairs and Digital Transformation of the Spanish Government and NextGenerationEU through the Quantum Spain project, and by CSIC Interdisciplinary Thematic Platform on Quantum Technologies (PTI-QTEP+). DFRE and JML are supported by Generalitat Valenciana (CIGRIS/2022/145 and ACIF/2021/219). MG is supported by CERN through the CERN QTI. Access to the IBM Quantum Services was obtained through the IBM Quantum Hub at CERN.

### v2

@@ V Conclusions
We have presented the first quantum computation of a total decay rate at second order in perturbative quantum field theory. Leveraging the loop-tree duality (LTD) framework, we have successfully combined loop and tree-level Feynman diagrams with a quantum algorithm on a quantum computer. This methodological advancement is significant from the high-energy physics perspective, as it allows us to integrate a real process with potential for quantum speedup. While we do not claim to have achieved quantum advantage in this work, our results lay the groundwork for future explorations in this direction. From the perspective of quantum computing, our study marks a noteworthy achievement. By solving a relatively complicated regression problem using a Quantum Neural Network (QNN) on a realistic dataset, we found a good compromise between trainability and expressibility, a common challenge in quantum neural networks. Most of the results presented are in agreement with the expected values within the uncertainty bands. This demonstrates the potential of quantum computing to address complex, real-world problems and highlights the importance of continuing to push the boundaries of what quantum technology can achieve.
@@ Acknowledgements.
This work is supported by the Spanish Government - Agencia Estatal de Investigación (MCIN/AEI/10.13039/501100011033) Grants No. PID2020-114473GB-I00, No. PID2023-146220NB-I00 and No. CEX2023-001292-S, and Generalitat Valenciana Grants No. PROMETEO/2021/071 and ASFAE/2022/009 (Planes Complementarios de I+D+i, NextGenerationEU). This work is also supported by the Ministry of Economic Affairs and Digital Transformation of the Spanish Government and NextGenerationEU through the Quantum Spain project and by the CSIC Interdisciplinary Thematic Platform on Quantum Technologies (PTI-QTEP+). DFRE and JML are supported by Generalitat Valenciana (CIGRIS/2022/145 and ACIF/2021/219). MG is supported by CERN through the CERN QTI. Access to the IBM Quantum Services was obtained through the IBM Quantum Hub at CERN.